# I Predict Therefore I Am: Is Next Token Prediction Enough to Learn Human-Interpretable Concepts from Data?

### v1

@@ 1 Introduction
Figure 1: An overview of the main contributions of this work. On the left, we illustrate the proposed latent variable model that represents concepts as latent variables $\mathbf{c}$, which are used to generate both the input $\mathbf{x}$ and output $y$ within a next-token prediction framework. Leveraging Bayes’ rule, the next-token prediction framework, and the diversity condition, we establish an identifiability result. Specifically, we demonstrate that the representations learned by LLMs approximately correspond to a linear transformation of the logarithm of the posterior distribution of latent variables conditioned on input tokens, i.e., $\mathbf{f}_{\mathbf{x}}(\mathbf{x})\approx\mathbf{A}\log{p(\mathbf{c}|\mathbf{x})}+\mathbf{k}$. This identifbaility results provide strong support for the linear representation hypothesis.
@@ 2 A Latent Variable Modal for Text Data
We present a novel latent variable model designed to capture the text data generation process, as depicted on the left in Figure 1. The proposed generative model can be expressed probabilistically as:
$p(\mathbf{x},y)=\sum_{\mathbf{c}}p(\mathbf{x}|\mathbf{c})p(y|\mathbf{c})p(\mathbf{c}),$ (1)
where latent variables $\mathbf{c}$ model latent human-interpretable concepts, and observed variables $\mathbf{x}$ and $y$ model words. We here distinguish observed variables into $\mathbf{x}$ and $y$ to align the input, i.e., context, and output token in the next-token prediction inference framework.
@@
Unlike many existing approaches (Jiang et al., 2024; Rajendran et al., 2024), we do not require the mapping from latent to observed variables to be invertible. This includes that which maps $\mathbf{c}$ to generate words $\mathbf{x}$, and that which maps $\mathbf{c}$ to generate word $y$. Allowing non-invertibility is a deliberate design choice driven by the fact that the mapping from latent to observed space is often highly complex and unknown. Imposing constraints on it may unnecessarily limit its flexibility. Moreover, in the context of text data, there are two key considerations. First, in natural language processing, the mapping from latent space to observed space often involves a many-to-one relationship. For example, different combinations of emotional concepts can lead to the same sentiment label. In sentiment analysis, the combination of ‘positive sentiment’ and ‘excitement’ might result in an observed outcome such as ‘This is amazing!’ (Miller, 1995). Second, some latent concepts may not be explicitly manifested in the observed sentence. For instance, a speaker’s intent, tone, or implicit connotations may influence how a sentence is constructed, yet remain unobservable in the surface-level text. This phenomenon is particularly prominent in pragmatics and discourse analysis, where unspoken contextual factors significantly shape meaning (Levinson, 1983). By not enforcing invertibility, our model can accommodate these unobservable latent factors, which are essential for robust and piratical identifiability analysis. This decision naturally implies that we may not achieve exact identifiability results as in previous works, due to inevitable information loss in the context of non-invertible mappings. Nevertheless, this non-invertibility still allows for the consideration of identifiability in an approximate sense, which will be discussed in Section 3.
3 Identifiabilty Analysis for Latent Concepts
@@
3.1 Next-token Prediction Approximately Identifies Latent Variable Posteriors
@@
where $Z(\mathbf{x})=\sum_{y^{\prime}}{\exp({\mathbf{f}_{\mathbf{x}}(\mathbf{x})^{T}\mathbf{f}_{y}(y^{\prime}))}}$. Equation (4) establishes an initial connection between the inference model (on the left) and the generative model (on the right). Intuitively, the first term on the left is a matrix-vector multiplication, while the right-hand side involves a logarithm of a summation over latent variables. To bridge this gap, we approximate the right-hand side to express it in a form similar to a matrix-vector multiplication, as follows:
$\mathbf{f}_{\mathbf{x}}(\mathbf{x})^{T}\mathbf{f}_{y}(y)-\log Z(\mathbf{x})\approx\sum_{\mathbf{c}}p(\mathbf{c}|y)\log p(\mathbf{c}|\mathbf{x})+\mathbf{b}_{y},$ (5)
where $\mathbf{b}_{y}$ does not depend on $\mathbf{c}$. If we assume that the mapping from latent space $\mathbf{c}$ to observed variables $\mathbf{x}$ approaches invertibility, the approximation in Eq. (5) becomes increasingly tight, which will be discussed in Section 3.2.
@@
Without additional assumptions, it is challenging to derive results beyond Eq. (5). To address this, we introduce the following assumption, referred to as the diversity condition. Formally, we assume that: there exist $\ell+1$ values of $y$, where $\ell$ is the dimension of $p(\mathbf{c})$ (equivalently $p(\mathbf{c}|y)$ or $p(\mathbf{c})|\mathbf{x}$), i.e., $y_{0}$, $y_{1}$, …, $y_{\ell}$, such that the matrix $\mathbf{L}=\big{(}p(\mathbf{c}|y=y_{1})-p(\mathbf{c}|y=y_{0}),...,p(\mathbf{c}|y=y_{\ell})-p(\mathbf{c}|y=y_{0})\big{)}$ of size $\ell\times\ell$ is invertible. This assumption was initially developed in the context of nonlinear ICA (independent component analysis) (Hyvarinen & Morioka, 2016; Hyvarinen et al., 2019; Khemakhem et al., 2020). Basically, it requires that the distribution $p(\mathbf{c}|y)$ exhibits sufficiently strong and diverse variation across different values of $y$. Similarly, we also assume that there exist $\ell+1$ values of $y$, so that the matrix $\mathbf{\hat{L}}=\big{(}\mathbf{f}_{y}(y=y_{1})-\mathbf{f}_{y}(y=y_{0}),...,\mathbf{f}_{y}(y=y_{\ell})-\mathbf{f}_{y}(y=y_{0})\big{)}$ of size $\ell\times\ell$ is also invertible. This assumption has likewise been employed in the context of identifiability analysis in inference space for LLMs (Roeder et al., 2021; Marconato et al., 2024). Under this diversity condition, we derive the following approximate identifiability result:
@@
Under the diversity condition above, the true latent variables $\mathbf{c}$ are related to the representations in LLMs, i,e, $\mathbf{f}_{\mathbf{x}}(\mathbf{x})$, which are learned through the next-token prediction framework, by the following relationship: $\mathbf{f}_{\mathbf{x}}(\mathbf{x})\approx\mathbf{A}\log{p(\mathbf{c}|\mathbf{x})}+\mathbf{k},$ where $\mathbf{A}=(\mathbf{\hat{L}}^{T})^{-1}\mathbf{L}$ is an invertible matrix, $\mathbf{k}$ denotes a constant vector.
Proof Sketch
The proof proceeds as follows. Expanding Eq. (5) at $\ell+1$ distinct points $y_{0},y_{1},\dots,y_{\ell}$ yields a system of $\ell+1$ equations. Subtracting the equation at $y_{0}$ from each of the equations at $y_{1},\dots,y_{\ell}$ provides $\ell$ equations. These $\ell$ equations can be combined into a single matrix equation involving $\mathbf{\hat{L}}$ and $\mathbf{L}$. By leveraging the invertibility of $\mathbf{\hat{L}}$, the result is obtained. Proof can be found in Appendix C.
@@
3.2 Approximation Tightness Analysis
We now examine the tightness of the approximate identifiability result. This approximation arises fundamentally from the invertibility of the mapping from the latent to observed space within the proposed latent variable model. In general, deriving a relative tight bound for such approximation requires specific properties of the mapping. However, since the mapping is unknown and highly complex in real applications, we refrain from imposing any constraints on it. Consequently, providing a reasonable bound for the approximation remains challenging. Nevertheless, technically speaking, such as in our proof process, the approximation arises from the transition from the right-hand side of Eq. (4) to that of Eq. (5). We can analyze this transition to gain valuable insights into the tightness of the approximation.
This transition is based on Jensen’s inequality applied to the concave logarithm function, as detailed in the following  [^1] See Eqs. (11) - (13) for further details.:
$\log\sum_{\mathbf{c}}p(\mathbf{c}|{y})\frac{p(y|\mathbf{c})p(\mathbf{c}|\mathbf{x})}{p(\mathbf{c}|{y})}\geq\sum_{\mathbf{c}}p(\mathbf{c}|{y})\log\frac{p(y|\mathbf{c})p(\mathbf{c}|\mathbf{x})}{p(\mathbf{c}|{y})}.$ (6)
We recommend the reader, who may be interested in the bound of Jensen’s inequality, to refer to Section F.
When the mapping from latent variables $\mathbf{c}$ to observed variables $\mathbf{x}$ approaches invertibility, the conditional distribution $p(\mathbf{c}|\mathbf{x})$ becomes highly concentrated, effectively behaving like a delta distribution centered at a unique point $\mathbf{c}^{*}$ corresponding to $\mathbf{x}$. This means that $p(\mathbf{c}|\mathbf{x})\approx 0$ for all $\mathbf{c}\neq\mathbf{c}^{*}$ and $p(\mathbf{c}^{*}|\mathbf{x})\approx 1$. Under this condition, the left-hand side of Eq. (6) simplifies as follows: $\log\sum_{\mathbf{c}}p(y=k|\mathbf{c})p(\mathbf{c}|\mathbf{x})=\log p(y=k|\mathbf{c}^{*})p(\mathbf{c}^{*}|\mathbf{x})$, because the summation is dominated by $\mathbf{c}^{*}$. Similarly, for the right-hand side of Eq. (6), $\sum_{\mathbf{c}}p(\mathbf{c}|{y})\log\frac{p(y|\mathbf{c})p(\mathbf{c}|\mathbf{x})}{p(\mathbf{c}|{y})}$ is also dominated by $\mathbf{c}^{*}$ due to the concentration of $p(\mathbf{c}|\mathbf{x})$. The dominance of $\mathbf{c}^{*}$ allows us to approximate $p(\mathbf{c}^{*}|{y})\approx 1$, since $\mathbf{c}^{*}$ now captures the overwhelming majority of the probability mass, rendering the contributions from other $\mathbf{c}$ values negligible. Therefore, $p(\mathbf{c}^{*}|{y})\approx 1$ becomes a valid approximation. As a result, the right-hand side is also approximately $\log p(y=k|\mathbf{c}^{*})p(\mathbf{c}^{*}|\mathbf{x})$. Thus, both sides approximately evaluate to the same value, ensuring a tight approximation.
In the analysis above, we assumed that the mapping from $\mathbf{c}$ to $\mathbf{x}$ approaches invertibility. However, when this mapping deviates from invertibility, the conditional distribution $p(\mathbf{c}|\mathbf{x})$ tends to become more uniform. This shift has important implications for the approximation. As the mapping becomes less invertible, the conditional distribution $p(\mathbf{c}|\mathbf{x})$ spreads out over a wider range of $\mathbf{c}$ values. In this scenario, the information that $\mathbf{x}$ provides about $\mathbf{c}$ diminishes, meaning $\mathbf{x}$ is less effective in pinpointing the specific value of $\mathbf{c}$. Consequently, the approximation above no longer focus narrowly around a particular $\mathbf{c}^{*}$, and instead the contributions from multiple values of $\mathbf{c}$ become more significant, which leads to a looser approximation.
This behavior is intuitive: when $p(\mathbf{c}|\mathbf{x})$ approaches a uniform distribution, we no longer expect to recover the exact value of $\mathbf{c}$. Without a clear focus on a particular value of $\mathbf{c}$, the approximation becomes less accurate because we cannot confidently identify the value of $\mathbf{c}$ that most contributes to the sum. Thus, the approximation’s precision deteriorates, reflecting the fact that the lack of invertibility reduces the effectiveness of the mapping and the ability to reliably estimate $\mathbf{c}$. Further empirical insights into the relationship between the degree of invertibility and latent variable recovery are presented in the left panel of Figure 2.
@@ 3.2 Approximation Tightness Analysis
Identifiability analysis is a fundamental challenge, especially in the causal representation learning community. While this community often assumes that the mapping from latent space to observed space is invertible (Brehmer et al., 2022; Von Kügelgen et al., 2021; Massidda et al., 2023; von Kügelgen et al., 2023; Ahuja et al., 2023; Seigal et al., 2022; Shen et al., 2022; Liu et al., 2022a), such an assumption may, in some real-world applications, oversimplify the inherently complex relationships between latent and observed variables. For example, in financial data, different combinations of latent factors (e.g., interest rates, geopolitical events) can lead to the same outcomes, making the mapping non-invertible. In healthcare, varied genetic and environmental factors can produce identical disease outcomes, while in recommendation systems, multiple latent features can result in the same interaction, all illustrating the non-invertibility of these mappings. These examples illustrate the non-invertibility of such mappings. We hope this work will inspire further research aimed at overcoming the limitations of invertibility assumptions in causal representation learning.
@@ 4.1 The Linear Representation Hypothesis
4.2 Understanding and Unifying the Hypothesis
@@
Given the relationship $\mathbf{f}_{\mathbf{x}}(\mathbf{x})\approx\mathbf{A}\log{p(\mathbf{c}|\mathbf{x})}+\mathbf{k}$ from Theorem 3.1, consider a counterfactual pair $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$ that only differs in the latent concept of interest, such that $c_{i}=0$ for $\mathbf{x}_{0}$ and $c_{i}=1$ for $\mathbf{x}_{1}$. The difference in the representations of $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$ is given by: $\mathbf{f}_{\mathbf{x}}(\mathbf{x}_{1})-\mathbf{f}_{\mathbf{x}}(\mathbf{x}_{0})\approx s\mathbf{A}_{:,i}$, where $s=\log p(c_{i}=1|\mathbf{x}_{1})-\log p(c_{i}=0|\mathbf{x}_{0})$ is a scaling factor.
Proof can be found in Appendix Section D.
@@
The corollary clarifies why the representation difference Rep(‘man’)-Rep(‘women’) can be approximated as Rep(‘king’)-Rep(‘queen’). Specifically, in both the ‘man’/‘woman’ and ‘king’/‘queen’ pairs, the primary difference stems from the male/female concept. This ensures that the representation differences in both pairs are aligned with the same column of $\mathbf{A}$. As a result, we can approximate Rep(‘men’)-Rep(‘women’) $\approx$ Rep(‘queen’)-Rep(‘king’), since both differences reflect the same underlying latent concept or the same column of $\mathbf{A}$.
@@
The corollary also supports the notion that a concept’s value can be adjusted by adding a corresponding steering vector, such as a specific column $\mathbf{A}_{:,i}$ in the matrix $\mathbf{A}$. Adding the steering vector effectively modifies the original representation to produce a new representation. For example, $\mathbf{\hat{f}}_{\mathbf{x}}=\mathbf{f}_{\mathbf{x}}(\mathbf{x})+\alpha\mathbf{A}_{:,i}=\mathbf{A}(\log p(\mathbf{c}|\mathbf{x})+\boldsymbol{\alpha})$, where $\alpha$ corresponds a weight as introduced (Wang et al., 2023; Turner et al., 2023), $\boldsymbol{\alpha}$ is a vector whose i-th element is $\alpha$ while all other elements are zero. Here, we observe that manipulating a concept using a steering vector is equivalent to altering the posterior $p(\mathbf{c}|\mathbf{x})$, which directly impacts the model’s output.
@@
Let $\mathbf{f}_{\mathbf{x}}(\mathbf{x})\approx\mathbf{A}\log{p(\mathbf{c}|\mathbf{x})}+\mathbf{k}$ as given in Theorem 3.1. For a data pair $(\mathbf{c},\mathbf{x})$, for a data pair $\mathbf{c}$ where $\mathbf{x}$ is generated by the latent $\mathbf{c}$ in the proposed latent variable model, the representations $\mathbf{f}_{\mathbf{x}}(\mathbf{x})$ satisfies that it can be classified linearly using a classifier with weights $\mathbf{W}$, and the weight matrix $\mathbf{W}$ is an invertible matrix related to $\mathbf{A}$.
Proof can be found in Appendix Section E.
The corollary supports the idea that concepts, such as the latent concept of English/French, can be predicted using a linear probe on the corresponding representations of the input. This is because the learned representations $\mathbf{f}_{\mathbf{x}}(\mathbf{x})$ in LLMs is a linear transformation of the log of the conditional distribution $p(\mathbf{c}|\mathbf{x})$. Such a transformation implies that a linear classifier is sufficient to align the model’s predicted distribution $p(\mathbf{c}|\mathbf{x})$ with the true distribution over classes $p(\mathbf{c})$. This alignment is achieved through the minimization of cross-entropy loss, ensuring effective prediction of the latent concepts.

### v2

@@ 1 Introduction
Figure 1: An overview of the main contributions of this work. On the left, we illustrate the proposed latent variable model that represents concepts as latent variables $\mathbf{c}$, which are used to generate both the input $\mathbf{x}$ and output $y$ within a next-token prediction framework. Leveraging Bayes’ rule, the next-token prediction framework, and the diversity condition, we establish an identifiability result. Specifically, we demonstrate that the representations learned by LLMs approximately correspond to a linear transformation of the logarithm of the posterior distribution of latent variables conditioned on input tokens, i.e., $\mathbf{f}_{\mathbf{x}}(\mathbf{x})=\mathbf{A}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}+o({\epsilon})$, where $o(\epsilon)$ represents a term that grows asymptotically smaller than $\epsilon$ as $\epsilon\to 0$. This identifiability results provide strong support for the linear representation hypothesis.
@@ 2 A Latent Variable Modal for Text Data
We present a novel latent variable model designed to capture the text data generation process, as depicted on the left in Figure 1. The proposed generative model can be expressed as:
$[\mathbf{x},y]=\mathbf{g}(\mathbf{c}),\quad\text{or probabilistically},\quad p(\mathbf{x},y)=\sum_{\mathbf{c}}p(\mathbf{x}|\mathbf{c})p(y|\mathbf{c})p(\mathbf{c}),$ (1)
where latent variables $\mathbf{c}$ model latent human-interpretable concepts, and observed variables $\mathbf{x}$ and $y$ model words generated by the deterministic mapping $\mathbf{g}$ on $\mathbf{c}$. We here distinguish observed variables into $\mathbf{x}$ and $y$ to align the input, i.e., context and output token in the next-token prediction inference framework, respectively.
@@
Unlike many existing approaches (Jiang et al., 2024; Rajendran et al., 2024), we do not require the mapping from latent to observed variables to be invertible. This includes that which maps $\mathbf{c}$ to generate words $\mathbf{x}$, and that which maps $\mathbf{c}$ to generate word $y$. Allowing non-invertibility is a deliberate design choice driven by the fact that the mapping from latent to observed space is often highly complex and unknown. Imposing constraints on it may unnecessarily limit its flexibility. Moreover, in the context of text data, there are two key considerations. First, in natural language processing, the mapping from latent space to observed space often involves a many-to-one relationship. For example, different combinations of emotional concepts can lead to the same sentiment label. In sentiment analysis, the combination of ‘positive sentiment’ and ‘excitement’ might result in an observed outcome such as ‘This is amazing!’ (Miller, 1995). Second, some latent concepts may not be explicitly manifested in the observed sentence. For instance, a speaker’s intent, tone, or implicit connotations may influence how a sentence is constructed, yet remain unobservable in the surface-level text. This phenomenon is particularly prominent in pragmatics and discourse analysis, where unspoken contextual factors significantly shape meaning (Levinson, 1983). By not enforcing invertibility, our model can accommodate these unobservable latent factors, which are essential for robust and piratical identifiability analysis. To this end, we define the approximate invertibility of the mapping $\mathbf{g}$ in Eq. (1) as follows:
Definition 2.1.
We define the degree of approximate invertibility of the mapping $\mathbf{g}$ in Eq. (1) by introducing an error term $\epsilon$, as follows:
$1-p(\mathbf{c}=\mathbf{c}^{*}|\mathbf{x},y)=\epsilon,$
where $10\leq\epsilon<$ 1, and $\mathbf{c}^{*}=\arg\max_{\mathbf{c}}p(\mathbf{c}|\mathbf{x},y)$ represents the dominant point.
Such relaxed condition naturally implies that we may not achieve exact identifiability results as in previous works, due to inevitable information loss in the context of non-invertible mappings. Nevertheless, this non-invertibility still allows for the consideration of identifiability in an approximate sense, which will be discussed in Section 3.
3 Identifiabilty Analysis for the Proposed Latent Variable Model
@@
where $Z(\mathbf{x})=\sum_{y^{\prime}}{\exp({\mathbf{f}_{\mathbf{x}}(\mathbf{x})^{T}\mathbf{f}_{y}(y^{\prime}))}}$. Equation (4) establishes an initial connection between the inference model (on the left) and the generative model (on the right). Intuitively, the first term on the left is a matrix-vector multiplication, while the right-hand side involves a logarithm of a summation over latent variables.
@@
Without additional assumptions, it is challenging to derive results beyond Eq. (4). To address this, we introduce the following assumption, referred to as the diversity condition. Formally, we assume that: there exist $\ell+1$ distinct values of $y$, where $\ell$ is the number of all possible values that $\mathbf{c}$ can take, i.e., $y_{0}$,…, $y_{k}$, …, $y_{\ell}$, such that the matrix $\mathbf{L}=\big{(}[p(\mathbf{c}=\mathbf{c}_{i}|y=y_{1})]_{\mathbf{c}_{i}}-[p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0})]_{\mathbf{c}_{i}},...,[p(\mathbf{c}=\mathbf{c}_{i}|y=y_{\ell})]_{\mathbf{c}_{i}}-[p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0})]_{\mathbf{c}_{i}}\big{)}$ of size $\ell\times\ell$ is invertible, where $[p(\mathbf{c}=\mathbf{c}_{i}|y=y_{1})]_{\mathbf{c}_{i}}$ denote the vector whose elements correspond to the probabilities of all possible values of $\mathbf{c}$, conditional on $y=y_{1}$. This assumption was initially developed in the context of nonlinear ICA (independent component analysis) (Hyvarinen & Morioka, 2016; Hyvarinen et al., 2019; Khemakhem et al., 2020). Basically, it requires that the distribution $p(\mathbf{c}|y)$ exhibits sufficiently strong and diverse variation across different values of $y$. Similarly, we also assume that there exist $\ell+1$ values of $y$, so that the matrix $\mathbf{\hat{L}}=\big{(}\mathbf{f}_{y}(y=y_{1})-\mathbf{f}_{y}(y=y_{0}),...,\mathbf{f}_{y}(y=y_{\ell})-\mathbf{f}_{y}(y=y_{0})\big{)}$ of size $\ell\times\ell$ is also invertible. This assumption has likewise been employed in the context of identifiability analysis in inference space for LLMs (Roeder et al., 2021; Marconato et al., 2024). Under this diversity condition, we derive the following approximate identifiability result:
@@
Under the diversity condition above, the true latent variables $\mathbf{c}$ are related to the representations in LLMs, i,e, $\mathbf{f}_{\mathbf{x}}(\mathbf{x})$, which are learned through the next-token prediction framework, by the following relationship:
$\mathbf{f}_{x}(\mathbf{x})=\mathbf{A}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}-(\mathbf{\hat{L}}^{T})^{-1}\mathbf{{h}}_{y}+(\mathbf{\hat{L}}^{T})^{-1}\mathbf{b}_{y},$ (5)
where ${\mathbf{h}_{y}}={[h_{y_{1}}-h_{y_{0}},...,h_{y_{\ell}}-h_{y_{0}}]}$ with $h_{y_{k}}[{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{k})}]^{T}_{\mathbf{c}_{i}}[\log{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{k},\mathbf{x})}]_{\mathbf{c}_{i}}$, and ${\mathbf{b}_{y}}={[b(y=y_{1})-b(y=y_{0}),...,b(y=y_{\ell})-b(y=y_{0})]}$ with $b(y=y_{k})=\mathbb{E}_{p(\mathbf{c}|y)}[\log{{p(y|\mathbf{c})}}]$, and $\mathbf{A}=(\mathbf{\hat{L}}^{T})^{-1}\mathbf{L}$. When $\epsilon=0$, $\mathbf{f}_{x}(\mathbf{x})=\mathbf{A}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}$. As $\epsilon\rightarrow 0$, $\mathbf{f}_{x}(\mathbf{x})\approx\mathbf{A}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}$.
@@
Identifiability analysis is a fundamental challenge, especially in the causal representation learning community. While this community often assumes that the mapping from latent space to observed space is invertible (Brehmer et al., 2022; Von Kügelgen et al., 2021; Massidda et al., 2023; von Kügelgen et al., 2023; Ahuja et al., 2023; Seigal et al., 2022; Shen et al., 2022; Liu et al., 2022), such an assumption may, in some real-world applications, oversimplify the inherently complex relationships between latent and observed variables. For example, in financial data, different combinations of latent factors (e.g., interest rates, geopolitical events) can lead to the same outcomes, making the mapping non-invertible. In healthcare, varied genetic and environmental factors can produce identical disease outcomes, while in recommendation systems, multiple latent features can result in the same interaction, all illustrating the non-invertibility of these mappings. These examples illustrate the non-invertibility of such mappings. We hope this work will inspire further research aimed at overcoming the limitations of invertibility assumptions in causal representation learning.
@@ 4.1 The Linear Representation Hypothesis
4.2 Understanding the Linear Representation Hypothesis
@@
Assume that Theorem 3.1 holds, i.e., $\mathbf{f}_{x}(\mathbf{x})\approx\mathbf{A}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}$. Define the degree of approximate invertibility of the mapping from $\mathbf{c}$ to $\mathbf{x}$ by $1-p(\mathbf{c}=\mathbf{c}_{\mathbf{x}}^{*}\mid\mathbf{x})=\epsilon_{\mathbf{x}}$, where $\mathbf{c}_{\mathbf{x}}^{*}$ denotes the dominant mode of $p(\mathbf{c}\mid\mathbf{x})$. Let $\mathbf{x}_{0}$ and $\mathbf{x}_{1}$ be two inputs that differ only in the value of the $i$th concept variable $c^{i}$, with $c^{i}=0$ for $\mathbf{x}_{0}$ and $c^{i}=1$ for $\mathbf{x}_{1}$, when $\epsilon_{\mathbf{x}}\rightarrow 0$, we have:
$\mathbf{f}_{\mathbf{x}}(\mathbf{x}_{1})-\mathbf{f}_{\mathbf{x}}(\mathbf{x}_{0})\approx s_{1}\mathbf{A}_{:,k_{1}}-s_{0}\mathbf{A}_{:,k_{0}},$
where $s_{1}\approx\log p(c^{i}=1|\mathbf{x}_{1})$, $s_{0}\approx\log p(c^{i}=0|\mathbf{x}_{0})$, and $k_{0}$ and $k_{1}$ be the indices of the concentrated mode $\mathbf{c}_{\mathbf{x}_{0}}^{*}$ for $\mathbf{x}_{0}$ and $\mathbf{c}_{\mathbf{x}_{1}}^{*}$ for $\mathbf{x}_{1}$, in the vector $\left[\log p(\mathbf{c}=\mathbf{c}_{j}\mid\mathbf{x})\right]_{\mathbf{c}_{j}}$, respectively.
@@
[…]
@@
The corollary also supports the notion that a concept’s value can be adjusted by adding a corresponding steering vector, such as $s_{1}\mathbf{A}_{:,k_{1}}-s_{0}\mathbf{A}_{:,k_{0}}$. Adding the steering vector effectively modifies the original representation to produce a new representation, i.e., $\mathbf{\hat{f}}_{\mathbf{x}}=\mathbf{f}_{\mathbf{x}}(\mathbf{x})+\alpha(s_{1}\mathbf{A}_{:,k_{1}}-s_{0}\mathbf{A}_{:,k_{0}})=\mathbf{A}([\log p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})]_{\mathbf{c}_{i}}+\boldsymbol{\alpha}_{s})$, where $\alpha$ corresponds a weight as introduced (Wang et al., 2023; Turner et al., 2023), $\boldsymbol{\alpha}_{s}$ is a vector whose $k_{0}$-th and $k_{1}$-th elements are $-\alpha*s_{0}$ and $\alpha*s_{1}$ respectively, while all other elements are zero. Thus, manipulating a concept via a steering vector is, in essence, equivalent to modifying the posterior distribution of the concept of interest, which directly impacts the model’s output.
@@
Assume that Theorem 3.1 holds, i.e., $\mathbf{f}_{x}(\mathbf{x})\approx\mathbf{A}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}$. Then, for a data pair $(\mathbf{c},\mathbf{x})$, where $\mathbf{x}$ is generated by the latents $\mathbf{c}$ in the proposed latent variable model, the representations $\mathbf{f}_{\mathbf{x}}(\mathbf{x})$ is linearly separable with respect to the latents $\mathbf{c}$, by a linear classifier with weight matrix $\mathbf{W}$, where $\mathbf{W}\mathbf{A}\approx\mathbf{I}$.
The corollary supports the idea that concepts, e.g., English/French, can be predicted using a linear probe on the corresponding representations of the input. This is because the learned representations $\mathbf{f}_{\mathbf{x}}(\mathbf{x})$ in LLMs is a linear transformation of the log of the conditional distribution $p(\mathbf{c}|\mathbf{x})$. Such a transformation implies that a linear classifier is sufficient to align the model’s predicted distribution $p(\mathbf{c}|\mathbf{x})$ with the true distribution over classes $p(\mathbf{c})$. This alignment is achieved through the minimization of cross-entropy loss, ensuring effective prediction of the latent concepts.